# Adaptive Width Neural Networks

### v4

@@ Abstract
For almost 70 years, researchers have primarily relied on hyper-parameter tuning to select the width of neural networks’ layers. This paper challenges the status quo by introducing an easy-to-use technique to learn an unbounded width of a neural network’s layer during training. The method jointly optimizes the width and the parameters of each layer via standard backpropagation. We apply the technique to a broad range of data domains such as tables, images, text, sequences, and graphs, showing how the width adapts to the task’s difficulty. A by product of our width learning approach is the easy truncation of the trained network at virtually zero cost, achieving a smooth trade-off between performance and compute resources. Alternatively, one can dynamically compress the network until performances do not degrade. In light of recent foundation models trained on large datasets, requiring billions of parameters and where hyper-parameter tuning is unfeasible due to huge training costs, our approach introduces a viable alternative for width learning.
@@ 1 Introduction
Since the construction of the Mark I Perceptron machine [48] the effective training of neural networks has remained an open research problem of great academic and practical value. The Mark I solved image recognition tasks by exploiting a layer of 512 fixed “association units” that in modern language correspond to the hidden units of a Multi-Layer Perceptron (MLP). MLPs possess universal approximation capabilities when assuming arbitrary width [10] and sigmoidal activations, and their convergence to good solutions was studied, for instance, in Rumelhart et al. [50] where the backpropagation algorithm was introduced as “simple, easy to implement on parallel hardware”, and improvable by other techniques such as momentum that preserve locality of weight updates.
@@
Yet, after almost 70 years of progress [34], the vast majority of neural networks, be they shallow or deep, still rely on a fixed choice of the number of neurons in their hidden layers. The width is typically treated as one of the many hyper-parameters that have to be carefully tuned whenever we approach a new task [62]. The tuning process has many names, such as model selection, hyper-parameter tuning, and cross-validation, and it is associated with non-negligible costs: Different architectural configurations are trained until one that performs best on a validation set is selected [41]. The configurations’ space grows exponentially in the number of layers, so practitioners often resort to shortcuts such as picking a specific number of hidden units for all layers, which greatly reduces the search space together with the chances of selecting a better architecture for the task. Other techniques to tune hyper-parameters include constructive approaches [17, 64], which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network [40], bi-level optimization [19], and neural architecture search [61], which often requires separate training runs for each configuration.
@@
The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language [9] and vision [68], for example, where parameters are in the order of billions to accommodate for a huge dataset. Training these models requires an amount of time, compute power, and energy that currently makes it unfeasible for most institutions to perform a thorough model selection and find good width parameters; the commonly accepted compromise is to stick to previously successful hyper-parameter choices. This may also explain why network pruning [6, 39], distillation [69] and quantization [39] techniques have recently been in the spotlight, as they trade-off hardware requirements and performance.
@@
This work introduces a simple and easy to use technique to learn the width of each neural network’s layer without imposing upper bounds (we refer to it as unbounded width). The width of each layer is dynamically adjusted during backpropagation [44], and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation. The technical strategy is to impose a soft ordering of hidden units by exploiting any monotonically decreasing function with unbounded support on natural numbers. With this, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks [61]. As a by-product, we can achieve a straightforward trade-off between parametrization and performance by deleting the last rows/columns of weight matrices, i.e., removing the “least important” neurons from the computation. Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which reduces the “jostling” effect where symmetric parametrizations compete when training starts [1].
@@
We test our method on MLPs for tabular data, a Convolutional Neural Network (CNN) [33] for images, a Transformer architecture [58] for text, a Recurrent Neural Network (RNN) for sequences, for and a Deep Graph Network (DGN) [38, 52] for graphs, to showcase a broad scope of applicability. Empirical results show, as may be intuitively expected, that the width adapts to the task’s difficulty with the same performance of a fixed-width baseline. We also investigate compressing of the network at training time while preserving accuracy, as well as a post-hoc truncation inducing a controlled trade-off at zero additional cost. Ablations suggest that the learned width is not influenced by the starting width (under bounded activations) nor by the batch size, advocating for a potentially large reduction of the hyper-parameter configuration space.
@@ 2 Related Work
Constructive methods dynamically learn the width of neural networks and are related in spirit to this work. The cascade correlation algorithm [17] alternates standard training with the creation of a new hidden unit minimizing the neural network’s residual error. Similarly, the firefly network descent [64] grows the width and depth of a network every $N$ training epochs via gradient descent on a dedicated loss. Yoon et al. [66] propose an ad-hoc algorithm for lifelong learning that grows the network by splitting and duplicating units to learn new tasks. Wu et al. [63] alternate training the network and then splitting existing neurons into offspring with equal weights. These works mostly focus on growing the neural network; Mitchell et al. [40] propose natural gradient-based heuristics to grow/shrink layers and hidden units of MLPs and CNNs. The main difference from our work is that we grow and shrink the network by simply computing the gradient of the loss, without relying on human-defined heuristics. The unbounded depth network of Nazaret and Blei [42], from which we draw inspiration, learns the number of layers of neural networks. Compared to that work, we focus our attention to the number of neurons, modifying the internals of the architecture rather than instantiating a multi-output one. Finally, we mention Bayesian nonparametrics approaches [43] that learn a potentially infinite number of clusters in an unsupervised fashion, as well as dynamic neural networks [24] that condition the architecture on input properties.
@@
Neural Architecture Search (NAS) is an automated process that designs neural networks for a given task [15, 61] and has been applied to different contexts [71, 35, 53]. Typically, neural network elements are added, removed, or modified based on validation performance [14, 60, 61], by means of reinforcement learning [70], evolutionary algorithms [47], and gradient-based approaches [35]. Typical NAS methods require enormous computational resources, sometimes reaching thousands of GPU days [70], due to the retraining of each new configuration. While recent advances on one-shot NAS models [8, 45, 2, 4, 55, 54] have drastically reduced the computational costs, they mostly focus on CNNs, assume a bounded search space, and do not learn the width. As such, NAS methods are complementary to our approach. Bi-level optimization algorithms have also been used for hyper-parameter tuning [19], where hyper-parameters are the variables of the outer objective and the model parameters those of the inner objective. The solution sets of the inner problem are usually not available in closed form, which has been partly addressed by repeated application of (stochastic) gradient descent [11, 36, 18]. These methods are restricted to continuous hyper-parameters’ optimization, and cannot therefore be applied to width optimization.
Finally, pruning [6] and distillation [29] are two methods that reduce the size of neural networks by trading-off performances; the former deletes neural connections [39] or entire neurons [57, 12], the latter trains a smaller network (student) to mimic a larger one (teacher) [22]. In particular, dynamic pruning techniques can compress the network at training time [23], by applying hard or soft masks [28]; for a comprehensive survey on pruning, please refer to [27]. Compared to most pruning approaches, our work can delete connections and reduce the model’s memory, but also grow it indefinitely; compared to distillation, we do not necessarily need a new training to compress the network. These techniques, however, can be easily combined with our approach.
@@ 3 Adaptive Width Learning
We now introduce Adaptive Width Neural Networks (AWNN), a probabilistic framework that maximizes a simple variational objective via backpropagation over a neural network’s parameters.
@@
[…]
@@
$p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)=\prod_{i=1}^{N}p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})=p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})p(\boldsymbol{\lambda})p(\boldsymbol{\theta})$ (2)
$p(\boldsymbol{\lambda})=\prod^{L}_{\ell=1}p(\lambda_{\ell})=\prod^{L}_{\ell=1}\mathcal{N}(\lambda_{\ell};\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell})\ \ \ \ \ \ p(\boldsymbol{\theta})=\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}p(\theta_{\ell n})\text{=}\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}\mathcal{N}(\theta_{\ell n};\mathbf{0},\text{diag}(\sigma^{\theta}_{\ell}))$ (3)
@@
Figure 1: (Left) The graphical model of AWNN, with dark observable random variables and white latent ones. (Middle) The distribution $f_{\ell}$ over hidden units’ importance at layer $\ell$ is parametrized by $\lambda_{\ell}$. The width of layer $\ell$ is chosen as the quantile function of the distribution $f_{\ell}$ evaluated at $k$ and denoted by $D_{\ell}$. (Right) The hidden units’ activations at layer $\ell$ are rescaled by their importance.
[…]
$\log p(Y|X)\geq\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right],$ (5)
where $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ is parametrized by learnable variational parameters. Before continuing, we define the truncated width $D_{\ell}$, that is the finite number of neurons at layer $\ell$, as the quantile function evaluated at $k$, with $k$ a hyper-parameter, of a distribution [^1] General functions are allowed if a threshold can be computed. $f_{\ell}$ with infinite support over $\mathbb{N}^{+}$ and parametrized by $\lambda_{\ell}$; Appendix A and B provide formal requirements about $f_{\ell}$. In other words, we find the integer such that the cumulative mass function of $f_{\ell}$ takes value $k$, and that integer $D_{\ell}$ is the truncated width of layer $\ell$. We implement $f_{\ell}$ as a discretized exponential distribution adhering to Def. A.2, following the discretization strategy of Roy [49]: For every natural $x$, the discretized distribution relies on the exponential’s cumulative distribution function:
$f_{\ell}(x;\lambda_{\ell})=(1-e^{\lambda_{\ell}(x+1)})-(1-e^{\lambda_{\ell}(x)}).$ (6)
@@
Then, we can factorize the variational distribution $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ into:
$q(\boldsymbol{\lambda},\boldsymbol{\theta})=q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda})$ $q(\boldsymbol{\lambda})=\prod_{\ell=1}^{L}q(\lambda_{\ell})=\prod_{\ell=1}^{L}\mathcal{N}(\lambda_{\ell};\nu_{\ell},1)$ (7)
$q(\boldsymbol{\theta}|\boldsymbol{\lambda})=\prod_{\ell=1}^{L}\prod_{n=1}^{D_{\ell}}q(\theta_{\ell n})\prod_{D_{\ell}+1}^{\infty}p(\theta_{\ell n})$ $q(\theta_{\ell n})=\mathcal{N}(\theta_{\ell n};\text{diag}(\rho_{\ell n}),\mathbf{I}).$ (8)
Here, $\nu_{\ell},\rho_{\ell n}$ are learnable variational parameters and, as before, we define $\boldsymbol{\rho}_{\ell}=\left\{\rho_{\ell n}\right\}_{n=1}^{D_{\ell}}$, $\boldsymbol{\rho}=\left\{\boldsymbol{\rho}_{\ell}\right\}_{\ell=1}^{L}$ and $\boldsymbol{\nu}=\left\{\nu_{\ell}\right\}_{\ell=1}^{L}$. Note that the set of variational parameters is finite as it depends on $D_{\ell}$.
By expanding Equation 5 using the above definitions and approximating the expectations at the first order, i.e., $\mathbb{E}_{q(\boldsymbol{\lambda})}[f(\boldsymbol{\lambda})]$=$f(\boldsymbol{\nu})$ and $\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[f(\boldsymbol{\theta})]=f(\boldsymbol{\rho})$ as in Nazaret and Blei [42], we obtain the final form of the ELBO [^2] The higher $k$ the better the ELBO approximation. (the full derivation is in Appendix C):
$\sum_{\ell}^{L}\log\frac{p(\nu_{\ell};\mu_{\ell}^{\lambda},\sigma_{\ell}^{\lambda})}{q(\nu_{\ell};\nu_{\ell})}+\sum_{\ell}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho_{\ell n};\sigma_{\ell}^{\theta})}{q(\rho_{\ell n};\rho_{\ell n})}+\sum_{i=1}^{N}\log p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i}),$ (9)
@@
In practice, the finite variational parameters $\boldsymbol{\nu},\boldsymbol{\rho}$ are those used to parametrize the neural network rather than sampling $\boldsymbol{\lambda},\boldsymbol{\theta}$, which enables easy optimization via backpropagation. Maximizing Equation (9) will update each variational parameter $\nu_{\ell}$, which in turn will change the value of $D_{\ell}$ during training. If $D_{\ell}$ increases we initialize new neurons and draw their weights from a standard normal distribution, otherwise we discard the weights of the extra neurons. When implementing mini-batch training, the predictive loss needs to be rescaled by $N/M$, where $M$ is the mini-batch size. From a Bayesian perspective, this is necessary as regularizers should weigh less if we have more data.
Compared to a fixed-width network with weight decay, we need to choose the priors’ values of $\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$, as well as initialize the learnable $\nu_{\ell}$. The latter can be initially set to same value across layers since they can freely adapt later, or it can be sampled from the prior $p(\lambda_{\ell})$. Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves [21, 5], so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWNN can converge to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values; iv) the more data, the less the priors will matter.

### v5

@@ Abstract
For almost 70 years, researchers have typically selected the width of neural networks’ layers either manually or through automated hyperparameter tuning methods such as grid search and, more recently, neural architecture search. This paper challenges the status quo by introducing an easy-to-use technique to learn an unbounded width of a neural network’s layer during training. The method jointly optimizes the width and the parameters of each layer via standard backpropagation. We apply the technique to a broad range of data domains such as tables, images, text, sequences, and graphs, showing how the width adapts to the task’s difficulty. A by product of our width learning approach is the easy truncation of the trained network at virtually zero cost, achieving a smooth trade-off between performance and compute resources. Alternatively, one can dynamically compress the network until performances do not degrade. In light of recent foundation models trained on large datasets, requiring billions of parameters and where hyper-parameter tuning is unfeasible due to huge training costs, our approach introduces a viable alternative for width learning.
@@ 1 Introduction
Since the construction of the Mark I Perceptron machine (Rosenblatt, 1958) the effective training of neural networks has remained an open research problem of great academic and practical value. The Mark I solved image recognition tasks with a layer of 512 fixed “association units” that in modern language are the hidden units of a Multi-Layer Perceptron (MLP). MLPs possess universal approximation capabilities when assuming arbitrary width (Cybenko, 1989) and sigmoidal activations, and their convergence to good solutions was studied, for instance, in Rumelhart et al. (1986) where the backpropagation algorithm was introduced as “simple, easy to implement on parallel hardware”, and improvable by other techniques such as momentum that preserve locality of weight updates.
@@
Yet, after almost 70 years of progress (LeCun et al., 2015), the vast majority of neural networks, be they shallow or deep, still rely on a fixed choice of the number of neurons in their hidden layers. The width is typically treated as one of the many hyper-parameters that have to be carefully tuned whenever we approach a new task (Wolpert, 1996). The tuning process has many names, such as model selection, hyper-parameter tuning, and cross-validation, and it is associated with non-negligible costs: Different architectural configurations are trained until one that performs best on a validation set is selected (Mitchell, 1997). The configurations’ space grows exponentially in the number of layers, so practitioners often resort to shortcuts such as picking a specific number of hidden units for all layers and trying a few values, which greatly reduces the search space together with the chances of selecting a better architecture for the task. Other techniques to tune hyper-parameters include constructive approaches (Fahlman & Lebiere, 1989; Wu et al., 2020), which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network (Mitchell et al., 2023), bi-level optimization (Franceschi et al., 2018), and neural architecture search (White et al., 2023), which often requires separate training runs for each configuration.
@@
The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language (Brown et al., 2020) and vision (Zhai et al., 2022), for example, where parameters are in the order of billions to accommodate for a huge dataset. Training these models requires an amount of time, compute power, and energy that currently makes it unfeasible for most institutions to perform a thorough model selection and find good width parameters; the commonly accepted compromise is to stick to previously successful hyper-parameter choices. This may also explain why network pruning (Blalock et al., 2020; Mishra et al., 2021), distillation (Zhang et al., 2019) and quantization (Mishra et al., 2021) techniques have recently been in the spotlight, as they trade-off hardware requirements and performance.
@@
This work introduces a simple and easy to use technique to learn the width of each neural network’s layer without imposing upper bounds (we refer to it as unbounded width). The width of each layer is dynamically adjusted during backpropagation (Paszke et al., 2017), and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation. The technical strategy is to impose a soft ordering of hidden units by exploiting any monotonically decreasing function with unbounded support on natural numbers. With this, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks (White et al., 2023). As a by-product, we can achieve a straightforward trade-off between parametrization and performance by deleting the last rows/columns of weight matrices, i.e., removing the “least important” neurons from the computation. Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which reduces the “jostling” effect where symmetric parametrizations compete when training starts (Barber, 2012).
@@
We test our method on MLPs for tabular data, a Convolutional Neural Network (CNN) (LeCun et al., 1989) for images, a Transformer (Vaswani et al., 2017) for text, a Recurrent Neural Network (RNN) for sequences, and a Deep Graph Network (DGN) (Micheli, 2009; Scarselli et al., 2009) for graphs, showcasing a broad scope of applicability as MLPs are ubiquitous in modern architectures. Empirical results show, as may be intuitively expected, that the width adapts to the task’s difficulty with performances comparable to fixed-width baseline. We also encourage compression of the network at training time while preserving accuracy, as well as a post-hoc truncation inducing a controlled trade-off at zero additional cost. Ablations suggest that the learned width is not influenced by the starting width (under bounded activations) nor by the batch size, advocating for a potentially large reduction of the hyper-parameter configuration space.
@@ 2 Related Work
Constructive methods dynamically learn the width of neural networks and are related in spirit to this work. The cascade correlation algorithm (Fahlman & Lebiere, 1989) alternates standard training with the creation of a new hidden unit minimizing the neural network’s residual error. Similarly, the firefly network descent (Wu et al., 2020) grows the width and depth of a network every $N$ training epochs via gradient descent on a dedicated loss. Yoon et al. (2018) propose an ad-hoc algorithm for lifelong learning that grows the network by splitting and duplicating units to learn new tasks. Wu et al. (2019) alternate training the network and then splitting existing neurons into offspring with equal weights. These works mostly focus on growing the neural network; Mitchell et al. (2023) propose natural gradient-based heuristics to grow/shrink layers and hidden units of MLPs and CNNs. The main difference from our work is that we grow and shrink the network by simply computing the gradient of the loss, without relying on human-defined heuristics. The unbounded depth network of Nazaret & Blei (2022), from which we draw inspiration, learns the number of layers of neural networks. Compared to that work, we focus our attention to the number of neurons, modifying the internals of the architecture rather than instantiating a multi-output one. In particular: i) learning the width requires a different formulation of the evidence lower-bound; and ii) the importance distribution needs to be monotonically decreasing, which does not encode the right inductive bias for the depth of a neural network. We also mention Bayesian nonparametric approaches (Orbanz & Teh, 2010) that learn a potentially infinite number of clusters in an unsupervised fashion, as well as dynamic neural networks (Han et al., 2021) that condition the architecture on input properties. Finally, the work of Caron et al. (2025) investigates, in a similar spirit to this work, how to introduce decreasing rescaling of the pre-activations of neurons in the context of very wide networks; the authors provide interesting global convergence results in the Neural Tangent Kernel (NTK) regime.
@@
Neural Architecture Search (NAS) is an automated process that designs neural networks for a given task (Elsken et al., 2019b; White et al., 2023) and has been applied to different contexts (Zoph et al., 2018; Liu et al., 2019; So et al., 2019). Typically, neural network elements are added, removed, or modified based on validation performance (Elsken et al., 2019a; White et al., 2021; 2023), by means of reinforcement learning (Zoph & Le, 2016), evolutionary algorithms (Real et al., 2019), and gradient-based approaches (Liu et al., 2019). Typical NAS methods require enormous computational resources, sometimes reaching thousands of GPU days (Zoph & Le, 2016), due to the retraining of each new configuration. While recent advances on one-shot NAS models (Brock et al., 2018; Pham et al., 2018; Bender et al., 2018; Berman et al., 2020; Su et al., 2021b; a) have drastically reduced the computational costs, they mostly focus on CNNs, assume a bounded search space, and do not learn the width. As such, NAS methods are complementary to our approach. Bi-level optimization algorithms have also been used for hyper-parameter tuning (Franceschi et al., 2018), where hyper-parameters are the variables of the outer objective and the model parameters those of the inner objective. The solution sets of the inner problem are usually not available in closed form, which has been partly addressed by repeated application of (stochastic) gradient descent (Domke, 2012; Maclaurin et al., 2015; Franceschi et al., 2017). These methods are restricted to continuous hyper-parameters’ optimization and cannot be applied to width optimization.
Finally, pruning (Blalock et al., 2020) and distillation (Hinton et al., 2015) are two methods that reduce the size of neural networks by trading-off performances; the former deletes neural connections (Mishra et al., 2021) or entire neurons (Valerio et al., 2022; Dufort-Labbé et al., 2024), the latter trains a smaller network (student) to mimic a larger one (teacher) (Gou et al., 2021). In particular, dynamic pruning techniques can compress the network at training time (Guo et al., 2016), by applying hard or soft masks (He et al., 2018); for a comprehensive survey on pruning, please refer to (He & Xiao, 2023). It is worth mentioning the pruning strategy of Wolinski et al. (2020), which promotes learning in some neurons and penalizes others by appropriate rescaling of a potentially (infinite-dimensional) input; this work shares conceptual similarities with the our approach, though it does not consider learning the rescaling factor as a proxy for the width. Compared to most pruning approaches, our work can delete connections and reduce the model’s memory, but also grow it indefinitely; compared to distillation, we do not necessarily need a new training to compress the network. These techniques can be easily combined with our approach, whose main goal is not compression but rather the automatic adaptation of a neural network’s width.
@@ 3 Adaptive Width Learning
We now introduce Adaptive Width Neural Networks (AWN), a probabilistic framework that maximizes a simple variational objective via backpropagation over a neural network’s parameters.
@@
We define $p(y|x)$ according to the graphical model of Figure 1 (left), to learn a neural network with unbounded width for each hidden layer $\ell$. To do so, we assume the existence of an infinite sequence of i.i.d. latent variables $\bm{\theta}_{\ell}=\left\{\theta_{\ell n}\right\}_{n=1}^{\infty}$, where $\theta_{\ell n}$ is a multivariate variable over the learnable weights of neuron $n$ at layer $\ell.$ However, since working with an infinite-width layer is not possible in practice, we also introduce a latent variable $\lambda_{\ell}$ that samples how many neurons to use for layer $\ell$. That is, it truncates an infinite width to a finite value so that we can feasibly perform inference with the neural network. For a neural network of $L$ layers, we define $\bm{\theta}=\left\{\bm{\theta}_{\ell}\right\}_{\ell=1}^{L}$ and $\bm{\lambda}=\left\{\lambda_{\ell}\right\}_{\ell=1}^{L}$, assuming independence across layers. Therefore, by marginalization one can write $p(Y|X)=\int p(Y,\bm{\lambda},\bm{\theta}|X)d\bm{\lambda}d\bm{\theta}$.
@@
$p(Y,\bm{\lambda},\bm{\theta}|X)=\prod_{i=1}^{N}p(y_{i},\bm{\lambda},\bm{\theta}|x_{i})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ p(y_{i},\bm{\lambda},\bm{\theta}|x_{i})=p(y_{i}|\bm{\lambda},\bm{\theta},x_{i})p(\bm{\lambda})p(\bm{\theta})$ (2)
$p(\bm{\lambda})=\prod^{L}_{\ell=1}p(\lambda_{\ell})=\prod^{L}_{\ell=1}\mathcal{N}(\lambda_{\ell};\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell})\ \ \ \ \ \ p(\bm{\theta})=\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}p(\theta_{\ell n})\text{=}\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}\mathcal{N}(\theta_{\ell n};\mathbf{0},\text{diag}(\sigma^{\theta}_{\ell}))$ (3)
@@
Figure 1: (Left) The graphical model of AWN, with dark observable random variables and white latent ones. (Middle) The distribution $f_{\ell}$ over hidden units’ importance at layer $\ell$ is parametrized by $\lambda_{\ell}$. The width of layer $\ell$ is chosen as the quantile function of the distribution $f_{\ell}$ evaluated at $k$ and denoted by $D_{\ell}$. (Right) The hidden units’ activations at layer $\ell$ are rescaled by their importance.
where $\sigma^{\theta}_{\ell},\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$ are hyper-parameters. For simplicity, we assumed a Gaussian prior for $\bm{\lambda}$, but please note that, when strict positivity is required, one should formally turn to Folded Normal priors. The neural network is parametrized by realizations $\bm{\lambda}\sim p(\bm{\lambda}),\bm{\theta}\sim\bm{\theta}$ – it relies on a finite number of neurons as detailed later and in Section 3.1 – and it outputs either class probabilities (classification) or the mean of a Gaussian distribution (regression) to parametrize $p(y_{i}|\bm{\lambda},\bm{\theta},x_{i})$ depending on the task. Maximizing Equation (1), however, requires computing the evidence $\int p(Y,\bm{\lambda},\bm{\theta}|X)d\bm{\lambda}d\bm{\theta}$, which is intractable. Therefore, we turn to mean-field variational inference (Jordan et al., 1999; Blei et al., 2017) to maximize an expected lower bound (ELBO) instead. This requires to define a variational distribution over the latent variables $q(\bm{\lambda},\bm{\theta})$ and re-phrase the objective as:
$\log p(Y|X)\geq\mathbb{E}_{q(\bm{\lambda},\bm{\theta})}\left[\log\frac{p(Y,\bm{\lambda},\bm{\theta}|X)}{q(\bm{\lambda},\bm{\theta})}\right],$ (5)
where $q(\bm{\lambda},\bm{\theta})$ is parametrized by learnable variational parameters. Before continuing, we define the truncated width $D_{\ell}$, that is the finite number of neurons at layer $\ell$, as the quantile function evaluated at $k$, with $k$ a hyper-parameter, of a distribution [^1] General functions are allowed if a threshold can be computed. $f_{\ell}$ with infinite support over $\mathbb{N}^{+}$ and parametrized by $\lambda_{\ell}$; Appendix A provides desirable properties of $f_{\ell}$ in a similar vein to Nazaret & Blei (2022). In other words, we find the integer such that the cumulative mass function of $f_{\ell}$ takes value $k$, and that integer $D_{\ell}$ is the truncated width of layer $\ell$. We implement $f_{\ell}$ as a discretized exponential distribution adhering to Def. A.2, following the discretization strategy of Roy (2003): For every $x\in\mathbb{N}^{+}$, the discretized distribution relies on the exponential’s cumulative distribution function:
$f_{\ell}(x;\lambda_{\ell})=(1-e^{-\lambda_{\ell}(x+1)})-(1-e^{-\lambda_{\ell}x}).$ (6)
@@
Then, we can factorize the variational distribution $q(\bm{\lambda},\bm{\theta})$ into:
$q(\bm{\lambda},\bm{\theta})=q(\bm{\lambda})q(\bm{\theta}|\bm{\lambda})$ $q(\bm{\lambda})=\prod_{\ell=1}^{L}q(\lambda_{\ell})=\prod_{\ell=1}^{L}\mathcal{N}(\lambda_{\ell};\nu_{\ell},1)$ (7)
$q(\bm{\theta}|\bm{\lambda})=\prod_{\ell=1}^{L}\prod_{n=1}^{D_{\ell}}q(\theta_{\ell n})\prod_{n^{\prime}=D_{\ell}+1}^{\infty}p(\theta_{\ell n^{\prime}})$ $q(\theta_{\ell n})=\mathcal{N}(\theta_{\ell n};\rho_{\ell n},\mathbf{I}).$ (8)
Here, $\nu_{\ell},\rho_{\ell n}$ are learnable variational parameters and, as before, we define $\bm{\rho}_{\ell}=\left\{\rho_{\ell n}\right\}_{n=1}^{D_{\ell}}$, $\bm{\rho}=\left\{\bm{\rho}_{\ell}\right\}_{\ell=1}^{L}$ and $\bm{\nu}=\left\{\nu_{\ell}\right\}_{\ell=1}^{L}$. Note that the set of variational parameters is finite as it depends on $D_{\ell}$.
By expanding Equation 5 using the above definitions and approximating the expectations at the first order, i.e., $\mathbb{E}_{q(\bm{\lambda})}[f(\bm{\lambda})]$$f(\bm{\nu})$ and $\mathbb{E}_{q(\bm{\theta}|\bm{\lambda})}[f(\bm{\theta})]=f(\bm{\rho})$ as in Nazaret & Blei (2022), we obtain the final form of the ELBO (the full derivation is in Appendix C):
$\sum_{\ell}^{L}\log\frac{p(\nu_{\ell};\mu_{\ell}^{\lambda},\sigma_{\ell}^{\lambda})}{q(\nu_{\ell};\nu_{\ell})}+\sum_{\ell}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho_{\ell n};\sigma_{\ell}^{\theta})}{q(\rho_{\ell n};\rho_{\ell n})}+\sum_{i=1}^{N}\log p(y_{i}|\bm{\lambda}\text{=}\bm{\nu},\bm{\theta}\text{=}\bm{\rho},x_{i}),$ (9)
@@
In practice, the finite variational parameters $\bm{\nu},\bm{\rho}$ are those used to parametrize the neural network rather than sampling $\bm{\lambda},\bm{\theta}$, which enables easy optimization via backpropagation. Maximizing Equation (9) will update each variational parameter $\nu_{\ell}$, which in turn will change the value of $D_{\ell}$ during training. If $D_{\ell}$ increases we initialize new neurons and draw their weights from a standard normal distribution, otherwise we discard the weights of the extra neurons. When implementing mini-batch training, the predictive loss needs to be rescaled by $N/M$, where $M$ is the mini-batch size. From a Bayesian perspective, this is necessary as regularizers should weigh less if we have more data.
Compared to a fixed-width network with weight decay, we need to choose the priors’ values of $\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$, as well as initialize the learnable $\nu_{\ell}$. The latter can be initially set to same value across layers since they can freely adapt later, or it can be sampled from the prior $p(\lambda_{\ell})$. Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves (Goel & Degroot, 1981; Bernardo & Smith, 2009), so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWN can converge to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values; iv) the more data, the less the priors will matter.